### v1

@@ 5 Numerical experiments
In this section, we demonstrate the efficiency of the proposed MPGDA algorithm by applying it to FSPCA problem (4) and SSC problem (7). All the experiments are implemented in MATLAB R2019a and conducted on a standard PC with 3.40GHz Intel(R) Core(TM) i5-7500 CPU and 16GB of RAM.
We first specify some implementation details of the proposed algorithm. Throughout the tests, we set $c_{1}=10^{-4}$, $\eta=0.1$, $l_{\min}=10^{-16}$, $l_{\max}=10^{16}$, $\tau_{1}=0.999$ and $\tau_{2}=0.9$. The parameters $\gamma_{0}$, $\xi_{0}>0$, $\theta>1$ and $\{T_{k}\}$ vary across the two application problems. Since $h$ is the $\ell_{1}$-norm in both problems, the semismooth Newton method developed in (chenProximalGradientMethod2020, , Section 4.2) is used to solve subproblem (17). We utilize the QR factorization as the retraction on the Stiefel manifold 2008Optimization .
@@ 5.1 Experiments on FSPCA
In this subsection, we compare the proposed MPGDA algorithm with the constraint relaxed majorization-minimization (CRMM) algorithm (babuFairPrincipalComponent2023, , Algorithm 3) for solving FSPCA problem (4). The CRMM algorithm first relaxes the semi-orthogonality constraint $X^{T}X=I$ into a linear matrix inequality constraint and then applies some majorization-minimization iteration procedure, which requires solving a quadratic program at each iteration.
We perform tests on synthetic datasets and a real dataset where the samples belong to two groups. The synthetic datasets are generated following similar settings to (2016SparseGeneralizedEigenvalueProblem, , Section 5.1). In detail, for the $i$-th group, we randomly generate $m_{i}=200$ samples with $d=40$ features according to Gaussian distribution with mean $\mu^{i}\in\mathbb{R}^{d}$ and covariance $\Sigma\in\mathbb{R}^{d\times d}$. We set $\mu_{j}^{1}=0,\ \mu_{j}^{2}=1/3$, for $j\in\{2,4,\cdots,200\}$, and $\mu_{j}^{i}=0$ otherwise. Meanwhile, the covariance matrix $\Sigma$ is chosen as a block diagonal matrix with five $8\times 8$ blocks, where the $(j,j^{\prime})$-th entry of each block is $0.8^{|j-j^{\prime}|}$. For the real-world data, we use the default credit dataset yeh2009creditdata and preprocess it as done in samadiPriceFairPCA . We set the regularization parameter $\mu=0.1$ in our tests. Furthermore, the CRMM algorithm is terminated if the relative error between the solutions obtained over two successive iterations is less than $10^{-6}$. For MPGDA algorithm, we set $\gamma_{0}=10^{-6}$, $\xi_{0}=4\sqrt{r}*10^{4}$, $\theta=1.5$, $T_{k}\equiv 15$, and terminate the algorithm if it satisfies:
$\mathcal{G}^{\beta_{k,0}}(x_{k},y_{k})<\varepsilon$ (25)
with $\varepsilon=10^{-6}$. Both compared algorithms start from the same random initial points on $\mathrm{St}(d,r)$ and the maximum iteration unmber of them are set as 1000.
We present the average results over 50 runs on different synthetic datasets and results for the default credit dataset in Table 1. We report the objective values (Obj), and CPU time in seconds (Time) and the number of iterations (Iter). It is observed that MPGDA algorithm achieves similar objective values to the CRMM algorithm while requiring much less computational time. Specially, our proposed algorithm is on average more than 20 times faster than the CRMM algorithm.
@@
CRMM MPGDA
Dataset r Obj. Time (s) Iter. Obj. Time (s) Iter.
Synthetic 2 $-9.795$ $0.269$ $201$9.808 0.007 167
3 $-14.380$ $0.344$ 155 -14.382 0.010 $192$
4 $-18.714$ $0.443$ 162 -18.745 0.012 $217$
5 $-22.774$ $0.209$ 83 -22.809 0.008 $197$
Credit 2 $-9.531$ $4.014$ 93 -9.742 0.133 $120$
3 $-10.557$ $15.692$ $271$10.802 0.104 91
4 $-11.637$ $8.977$ $142$11.923 0.010 5
5 $-12.572$ $20.058$ $277$12.724 0.012 4
@@ 5.2 Experiments on SSC
In this subsection, we compare the proposed MPGDA method with the RADA-RGD and RADA-PGD algorithms in xuRiemannianAlternatingDescent2024 for solving the SSC problem. Note that in xuRiemannianAlternatingDescent2024 , the original SSC problem (5) is reformulated into the following smooth minimax problem on the Grassmann manifold so that the RADA-PGD and RADA-RGD algorithms can be directly applied:
@@
Here, $\mathrm{Gr}(N,p)=\{XX^{\top}\in\mathbb{R}^{N\times N}\mid X\in\mathrm{St}(N,p)\}$ denotes the Grassmann manifold AGrassmannmanifoldhandbook2024A . The codes of these two algorithms for solving (26) are downloaded from: https://github.com/XuMeng00124/RADAopt.
In the tests, we generate synthetic datasets as in 2016ConvexSparseSpectralClustering , where the data points $\{a_{i}\}_{i=1}^{N}$ are independently drawn from the standard Gaussian distribution and $W_{ij}=|\langle a_{i},a_{j}\rangle|$. For the two RADA algorithms, we adopt the parameters as suggested in (xuRiemannianAlternatingDescent2024, , Section 6.3). Specifically, they are terminated once an $\varepsilon$-game-stationary point of problem (26) is found with $\varepsilon=10^{-4}$ and the initial point is set as $X_{0}X_{0}^{\top}$, where $X_{0}$ consists of $p$ eigenvectors associated with the $p$ smallest eigenvalues of $L$. For MPGDA algorithm, we set $\gamma_{0}=10^{-5}$, $\xi_{0}=\sqrt{p}N^{2}$, $\theta=2$ and $T_{k}\equiv 3$. Moreover, we choose the aforementioned $X_{0}$ as the initial point of MPGDA algorithm and terminate the method if it satisfies (25) with $\varepsilon=10^{-4}$. The maximum number of iterations for all the compared algorithms is set as 1000.
@@
Config Para $(p/\mu)$ RADA-PGD RADA-RGD MPGDA
@@
$\mu=0.1$ $N=200$ 2 2.187 $0.513$ $265$ 2.187 $0.796$ $153$ 2.187 0.114 65
4 $4.374$ $0.530$ $259$ 4.373 $0.812$ $157$ $4.374$ 0.116 64
6 6.560 $0.521$ $224$ 6.560 $0.855$ $164$ 6.560 0.117 63
8 8.747 $0.580$ $198$ 8.747 $0.861$ $166$ 8.747 0.129 64
10 10.934 $0.621$ $180$ 10.934 $0.995$ $167$ 10.934 0.129 65
$p=5$ $N=200$ 0.1 5.467 $0.535$ $244$ 5.467 $0.868$ $164$ 5.467 0.122 63
0.2 5.967 $0.556$ $263$ 5.967 $0.934$ $178$ 5.967 0.133 71
0.5 7.467 $0.593$ $286$ 7.467 $1.071$ $202$ 7.467 0.174 86
1.0 9.967 $0.616$ $300$ 9.967 $1.188$ $228$ 9.967 0.205 95
The computational results averaged over 50 synthetic data sets with $N=200$ data points are summarized in Table 2, where we report the obtained objective function value of problem (5) (Obj.), the CPU time in seconds (Time), and number of outer iterations (Iter.). We observe that all the compared algorithms attain comparable objective function value for different number of groups and regularization parameters $\mu$. In addition, the proposed MPGDA method consistently outperform the two RADA algorithms in terms of CPU time, which again validates its computational efficiency.

### v2

@@ 6 Numerical experiments
In this section, we illustrate the proposed algorithms via three numerical examples. We begin with an analytic example of problem (1), where the objective function is not linear with respect to $y$. This example is used to illustrate the differences in behavior between the MPGDA-PA algorithm and the MPGDA-PGA algorithm. Then, we examine the performance of the MPGDA-PA algorithm for solving the FSPCA problem (5) and the SSC problem (8) respectively. Since these two problems are nonconvex-linear, the two proposed algorithms have comparable computational cost at each iteration, but the MPGDA-PA algorithm enjoys better outer iteration complexity. Hence, it is more reasonable to apply MPGDA-PA rather than MPGDA-PGA algorithm for solving problems (5) and (8). All experiments are implemented in MATLAB R2025a and conducted on a standard PC with 3.40GHz Intel(R) Core(TM) i5-7500 CPU and 24GB of RAM.
We first specify some implementation details of the proposed MPGDA-PA algorithm. Throughout the tests, we set $c_{1}=10^{-4}$, $\eta=0.1$, $l_{\min}=10^{-16}$, $l_{\max}=10^{16}$, $\delta_{0}=10^{10}$, $\tau_{1}=0.999$, and $\tau_{2}=0.9$. The parameters $\gamma_{0}$, $\xi_{0}>0$, $\theta>1$ and $\{T_{k}\}$ vary across the three examples. Since $h$ is the $\ell_{1}$-norm in both problems, the semismooth Newton method developed in (chenProximalGradientMethod2020, Section 4.2) is used to solve subproblem (23). We utilize the QR factorization as the retraction on the Stiefel manifold 2008Optimization.
6.1 Experiments on an Analytic Minimax Problem
Consider the following nonconvex-concave minimax problem on $\mathbb{R}^{2}\times\mathbb{R}$:
$\min_{x_{1}^{2}+x_{2}^{2}=1}\max_{0.3\leq y\leq 1}-0.01x_{1}^{3}y-y\ln(y),$ (60)
which is a special case of problem (1) with $\mathcal{M}$ being the Stiefel manifold $\mathrm{St}(2,1)$, $S$ being the interval $[0.3,1]$, and $f(x,y)=-0.01x_{1}^{3}y-y\ln(y)$. A direct verification shows that $(1,0)$ is the unique minimizer to the value function of (60) and $(x^{*},y^{*}):=(1,0,e^{-1.01})$ is a game-stationary point. Now we apply the two proposed algorithms to solve (60). For the MPGDA-PA algorithm, we set $\gamma_{0}=0.005$, $\xi_{0}=1$, $\theta=1.5$, and $T_{k}=1$. We use the bisection method to solve the corresponding $y$-subproblem (21), which is implemented via the fminbnd command in MATLAB. For MPGDA-PGA, we set $c_{1}=10^{-4}$, $\hat{l}_{\min}=10^{-16}$, $\hat{l}_{\max}=10^{8}$, $\rho=0.2$, $\kappa=10^{16}$, and $\eta=0.5$. Both algorithms are initialized from the same starting point $(0.8,0.6,0.3)\in\mathrm{St}(2,1)\times[0.3,1]$. We use the Euclidean distance between the iterates $(x_{k},y_{k})$ and the optimal point $(x^{*},y^{*})$, denoted by $D_{k}$, to measure the accuracy of the iterates generated by them.
Table 1: Comparison of iteration numbers and CPU time for reaching different levels of distance
Level Iter. Time
MPGDA-PA MPGDA-PGA MPGDA-PA MPGDA-PGA
$1\times 10^{-2}$ 17 918 0.038659 0.023850
$1\times 10^{-3}$ 19 2100 0.044875 0.054841
$3\times 10^{-4}$ 21 2767 0.058287 0.073542
$2\times 10^{-4}$ 38 3067 0.114491 0.084889
$1.5\times 10^{-4}$ 88 3455 0.167304 0.118553
Table 1 presents the first outer iteration number ( Iter.) required by the proposed algorithms to reach different levels of distance, and the corresponding CPU time in seconds (Time). Moreover, Figures (1(a)) and (1(b)) show $D_{k}$ (in logarithmic scale) versus iteration number $k$ and $D_{k}$ (in logarithmic scale) versus CPU time in seconds respectively. We observe that the iterates generated by both proposed algorithms converge to the optimal solution $(x^{*},y^{*})$. Although the MPGDA-PA algorithm requires to solve a nonlinear and strongly concave subproblem with respect to $y$ in each iteration, which results in a higher computational cost per iteration, it converges in substantially fewer outer iterations. In contrast, MPGDA-PGA performs simpler iterations at a much lower cost per iteration, but requires much more iterations to achieve comparable accuracy. Consequently, the overall CPU times of both algorithms are similar across different levels of accuracy, demonstrating a consistent trade-off between per-iteration cost and convergence speed for solving problem (60).
(a) $D_{k}$ (in log scale) vs Iters
(b) $D_{k}$ (in log scale) vs Time
Figure 1: Plots of $D_{k}$ (in log scale) for the proposed algorithms.
@@ 6.2 Experiments on FSPCA
In this subsection, we compare the proposed MPGDA-PA algorithm with the constraint relaxed majorization-minimization (CRMM) algorithm (babuFairPrincipalComponent2023, Algorithm 3) for solving FSPCA problem (5). The CRMM algorithm first relaxes the semi-orthogonality constraint $X^{T}X=I$ into a linear matrix inequality constraint and then applies some majorization-minimization iteration procedure, which requires solving a quadratic program at each iteration.
We perform tests on synthetic datasets and a real dataset where the samples belong to two groups. The synthetic datasets are generated following similar settings to (2016SparseGeneralizedEigenvalueProblem, Section 5.1). In detail, for the $i$-th group, we randomly generate $m_{i}=200$ samples with $d=40$ features according to the Gaussian distribution with mean $\mu^{i}\in\mathbb{R}^{d}$ and covariance $\Sigma\in\mathbb{R}^{d\times d}$. We set $\mu_{j}^{1}=0,\ \mu_{j}^{2}=1/3$, for $j\in\{2,4,\cdots,200\}$, and $\mu_{j}^{i}=0$ otherwise. Meanwhile, the covariance matrix $\Sigma$ is chosen as a block diagonal matrix with five $8\times 8$ blocks, where the $(j,j^{\prime})$-th entry of each block is $0.8^{|j-j^{\prime}|}$. For the real-world data, we use the default credit dataset yeh2009creditdata and preprocess it as done in samadiPriceFairPCA. We set the regularization parameter $\mu=0.1$ in our tests. Furthermore, the CRMM algorithm is terminated if the relative error between the solutions obtained over two successive iterations is less than $10^{-6}$. For the MPGDA-PA algorithm, we set $\gamma_{0}=10^{-6}$, $\xi_{0}=4\sqrt{r}*10^{4}$, $\theta=1.5$, $T_{k}\equiv 15$, and we terminate the algorithm if it satisfies:
$\mathcal{G}^{\beta_{k,0}}(x_{k},y_{k})<\varepsilon\text{ with }\varepsilon=10^{-6}.$ (61)
Both compared algorithms start from the same random initial points on $\mathrm{St}(d,r)$ and the maximum iteration number of them are set as 1000.
We present the average results over 50 runs on different synthetic datasets and results for the default credit dataset in Table 2. We report the objective value (Obj), CPU time in seconds (Time) and the number of iterations (Iter). It is observed that the MPGDA-PA algorithm achieves similar objective values to the CRMM algorithm while requiring much less computational time. Specifically, our proposed algorithm is on average more than 20 times faster than the CRMM algorithm.
@@
CRMM MPGDA-PA
Dataset r Obj. Time Iter. Obj. Time Iter.
Synthetic 2 9.802-9.802 $0.122$ $204$ 9.820-9.820 0.005 $167$
3 14.399-14.399 $0.131$ $165$ 14.401-14.401 0.007 $189$
4 18.755-18.755 $0.175$ $148$ 18.776-18.776 0.008 $216$
5 22.840-22.840 $0.089$ $79$ 22.867-22.867 0.007 $216$
Credit 2 9.531-9.531 $4.305$ $93$ 9.742-9.742 0.276 $124$
3 10.557-10.557 $14.599$ $272$ 10.802-10.802 0.138 $91$
4 11.637-11.637 $6.478$ $143$ 11.923-11.923 0.013 55
5 12.572-12.572 $14.332$ $274$ 12.724-12.724 0.013 44
@@ 6.3 Experiments on SSC
In this subsection, we compare the proposed MPGDA-PA method with the RADA-RGD and RADA-PGD algorithms in xuRiemannianAlternatingDescent2024 for solving the SSC problem. Note that in xuRiemannianAlternatingDescent2024, the original SSC problem (6) is reformulated into the following smooth minimax problem on the Grassmann manifold so that the RADA-PGD and RADA-RGD algorithms can be directly applied:
@@
Here, $\mathrm{Gr}(N,p)=\{XX^{\top}\in\mathbb{R}^{N\times N}\mid X\in\mathrm{St}(N,p)\}$ denotes the Grassmann manifold AGrassmannmanifoldhandbook2024A. The code of these two algorithms for solving (62) is downloaded from: https://github.com/XuMeng00124/RADAopt. In the tests, we generate synthetic datasets as in 2016ConvexSparseSpectralClustering, where the data points $\{a_{i}\}_{i=1}^{N}$ are independently drawn from the standard Gaussian distribution and $W_{ij}=|\langle a_{i},a_{j}\rangle|$. For the two RADA algorithms, we adopt the parameters as suggested in (xuRiemannianAlternatingDescent2024, Section 6.3) except the number of inner iterations ($T_{k}$). Specifically, they are terminated once an $\varepsilon$-game-stationary point of problem (62) is found with $\varepsilon=10^{-4}$ and the initial point is set as $X_{0}X_{0}^{\top}$, where $X_{0}$ consists of $p$ eigenvectors associated with the $p$ smallest eigenvalues of $L$. For the MPGDA-PA algorithm, we set $\gamma_{0}=10^{-5}$, $\xi_{0}=\sqrt{p}N^{2}$, and $\theta=2$. Moreover, we choose the aforementioned $X_{0}$ as the initial point of the MPGDA-PA algorithm and terminate the method if it satisfies (61) with $\varepsilon=10^{-4}$. The maximum number of iterations for all the compared algorithms is set as 1000.
In the experiments, we generate 50 synthetic datasets with $N=200$ data points. To fine tune the parameters $T_{k}$ involved, we test all the compared algorithms with $T_{k}\in\{1,2,3,4,5\}$ for solving the SSC problem with various choices of $p$ and $\mu$. The results reveal that the RADA-PGD, RADA-RGD and MPGDA-PA algorithms consistently achieve optimal performance with $T_{k}=1$, $2$, and $3$, respectively. Due to the limitation of space, we only report one representative result in Table 3, which shows the performance evaluation for different choices of $T_{k}$ on the SSC problem with $(p,\mu)=(5,1.0)$. Also, the computational results for the compared algorithms with their respective optimal $T_{k}$ are summarized in Table 4. Both tables present the objective function value of problem (6) (Obj.), the CPU time in seconds (Time), and number of outer iterations (Iter.) averaged over the 50 datasets. We observe that all the compared algorithms attain comparable objective function values for different numbers of groups and regularization parameters $\mu$. In addition, the proposed MPGDA-PA method consistently outperforms the two RADA algorithms in terms of CPU time, which again validates its computational efficiency. The superior performance of the proposed MPGDA-PA algorithm may be attributed to the fact that it is able to directly address the nonsmooth $\ell_{1}$ norm in the SSC problem, while the RADA algorithms can only tackle the converted smooth minimax formulation of SSC.
Table 3: Performance evaluation for different choices of $T_{k}$ on the SSC problem with $(p,\mu)=(5,1.0)$
RADA-PGD RADA-RGD MPGDA-PA
$T_{k}$ Obj. Time Iter. Obj. Time Iter. Obj. Time Iter.
1 $9.971$ 0.578 $300$ $17.590$ $1.787$ $1000$ $22.320$ $2.063$ $1000$
2 $9.971$ $1.148$ $300$ $10.070$ 0.873 $238$ $10.029$ $2.080$ $848$
3 $9.971$ $1.629$ $297$ $9.967$ $1.170$ $217$ $9.968$ 0.256 $97$
4 $9.971$ $2.198$ $301$ $9.967$ $1.830$ $254$ $9.968$ 0.298 $88$
5 $9.971$ $2.516$ $300$ $64.748$ $2.742$ $289$ $9.970$ 0.311 $84$
@@
Parameters
($p$, $\mu$)
RADA-PGD RADA-RGD MPGDA-PA
@@
(22, $0.1$) $2.187$ $0.408$ $265$ $2.187$ $0.451$ $154$ $2.187$ 0.106 $66$
(44, $0.1$) $4.374$ $0.410$ $259$ $4.383$ $0.517$ $180$ $4.373$ 0.101 $64$
(66, $0.1$) $6.560$ $0.400$ $224$ $6.569$ $0.643$ $223$ $6.560$ 0.103 $63$
(88, $0.1$) $8.747$ $0.431$ $198$ $8.753$ $0.765$ $264$ $8.747$ 0.109 $66$
($10$, $0.1$) $10.934$ $0.474$ $180$ $10.956$ $0.943$ $277$ $10.934$ 0.113 $64$
(55, $0.1$) $5.467$ $0.416$ $244$ $5.494$ $0.719$ $240$ $5.467$ 0.110 $65$
(55, $0.2$) $5.967$ $0.437$ $263$ $5.977$ $0.636$ $217$ $5.967$ 0.117 $71$
(55, $0.5$) $7.469$ $0.467$ $286$ $7.489$ $0.672$ $229$ $7.467$ 0.155 $86$
(55, $1.0$) $9.971$ $0.490$ $300$ $10.070$ $0.702$ $238$ $9.968$ 0.189 $97$
Funding Qia Li is supported by the National Natural Science Foundation of China under grant 12471098, and by the Guangdong Province Key Laboratory of Computational National Science at Sun Yat-sen University Grant 2020B1212060032.
Data Availability The default credit dataset used in Section 6.2 was downloaded from the website https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients.
Declarations
Conflict of interest  The authors declare that they have no conflict of interest.